OpenStudy (anonymous):

Factor by grouping. x^3+3x^2-8x-24

OpenStudy (anonymous):

Can I cheat ?

myininaya (myininaya):

\[(x^3+3x^2)+(-8x-24)=x^2(x+3)+(-8)(x+3)=(x+3)(x^2-8)\]

OpenStudy (anonymous):

this is a job for .... oh hello!

OpenStudy (anonymous):

oh wow you did that fast

OpenStudy (anonymous):

she is a factor by grouping machine

myininaya (myininaya):

lol